# Ontology-driven Prompt Tuning for LLM-based Task and Motion Planning

### v1

@@ VI Conclusion
This work introduced a novel ontology-driven prompt-tuning framework to enhance task and motion planning with large language models. Integrating knowledge-based reasoning and ontology-driven environment state descriptions, the framework dynamically refines user prompts to generate semantically accurate and context-aware symbolic plans. Unlike static, template-based approaches, it addresses flaws such as incorrect temporal goal ordering, and adapts effectively to dynamic environments. Validation in both simulation and real-world scenarios demonstrated significant improvements in planning accuracy and adaptability, highlighting the potential of combining LLMs with ontological reasoning for advanced robotic planning in complex tasks and dynamic contexts.

### v2

@@ VI Conclusions
This work introduced a novel Onto-LLM-TAMP framework to enhance task and motion planning with LLMs. Integrating knowledge-based reasoning and ontology-driven environment state descriptions, the framework dynamically refines user prompts to generate semantically accurate and context-aware symbolic plans. Unlike static, template-based approaches, it addresses flaws such as incorrect temporal goal ordering and adapts effectively to dynamic environments. Validation in both simulation and real-world scenarios demonstrated significant improvements in planning accuracy and adaptability, highlighting the potential of combining LLMs with ontological reasoning for advanced robotic planning in complex tasks and dynamic contexts.
Acknowledgement
This work has been supported by the Khalifa University of Science and Technology under Award No. RC1-2018-KUCARS, and by the European Commission’s Horizon Europe Framework Programme with the project IntelliMan (AI-Powered Manipulation System for Advanced Robotic Service, Manufacturing and Prosthetics) under Grant Agreement 101070136.
Declaration
During the preparation of this work, the author (s) used ChatGPT and Gemini to improve language and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and takes (s) full responsibility for the content of the publication.